# Hard-Constrained Neural Networks with Universal Approximation Guarantees

### v1

@@ 5.3 Optimizing Safe Control Policy
NN 350.77 (0.45) 151.56 (0.45) 110.62 (0.38) 0.30 (0.01)
NN+Proj 1488.46 (440.51) 0.00 (0.00) 0.00 (0.00) 3.98 (0.06)
Soft 412.37 (0.80) 12.86 (0.84) 7.33 (0.51) 0.28 (0.06)
Soft+Proj 3296.87 (5622.82) 0.00 (0.00) 0.00 (0.00) 4.35 (2.53)
DC3 429.84 (34.73) 15.55 (4.06) 10.47 (4.15) 38.63 (0.13)
HardNet-Aff 461.97 (14.35) 0.00 (0.00) 0.00 (0.00) 3.05 (0.69)
Figure 3: Simulated trajectories from a random initial state. The cost of each trajectory is shown in parentheses. HardNet-Aff avoids the obstacles while obtaining a low cost value. Even though the soft-constrained method and DC3 appear to avoid obstacles and achieve smaller costs than the other collision-free trajectories, they violate the safety constraints (which are more conservative than hitting the obstacles).
In this experiment, we apply HardNet-Aff to enforce safety constraints in control systems. Consider a control-affine system with its dynamics:
$\dot{x}(t)=f(x(t))+g(x(t))u,$ (11)
where $x(t)\in\mathbb{R}^{n_{\text{in}}}$ is the system state at time $t$ and $u\in\mathbb{R}^{n_{\text{out}}}$ is the control input. For safety reasons (e.g., avoiding obstables), the system requires $x(t)\in\mathcal{X}_{\text{safe}}\subset\mathbb{R}^{n_{\text{in}}}$ for all $t$. We translate this safety condition into a state-dependent affine constraint on the control input using a control barrier function (CBF) $h:\mathbb{R}^{n_{\text{in}}}\rightarrow\mathbb{R}$ (Ames et al., 2019). Suppose its super-level set $\{x\in\mathbb{R}^{n_{\text{in}}}|h(x)\geq 0\}\subset\mathcal{X}_{\text{safe}}$ and $h(x(0))\geq 0$. Then, we can ensure $h(x(t))\geq 0\;\forall t\geq 0$ by guaranteeing
@@
We consider controlling a unicycle system to minimize the cost over trajectories while avoiding collisions with two elliptical obstacles, each presented with a CBF (see Appendix A.5 for details). Given a nominal controller $\pi_{\text{nom}}:\mathbb{R}^{n_{\text{in}}}\rightarrow\mathbb{R}^{n_{\text{out}}}$ designed without obstacle considerations, a conventional approach to find a safe controller is to solve a quadratic program:
@@
at each $x(t)$. The downside of this method is that the controller cannot optimize a cost/reward over the trajectories as it only attempts to remain close to the nominal controller. Instead, we can do so by training neural network policies $\pi_{\theta}(x):=\pi_{\text{nom}}(x)+f_{\theta}(x)$ with neural networks $f_{\theta}$ by minimizing the costs of rolled-out trajectories from randomly sampled initial states. As shown in Fig. 3 and Table 4, HardNet-Aff consistently generates safe trajectories with low costs.
5.4 Airborne Collision Avoidance System (ACAS)
In this example, we consider learning an airborne collision avoidance system called the Horizontal Collision Avoidance System (HCAS), which is a variant of the popular system ACAS Xu used by Katz et al. (2017), for which the labeled training dataset is publicly available (Julian & Kochenderfer, 2019). The goal of the HCAS system is to recommend horizontal maneuvers—clear of conflict, weak left, strong left, weak right, strong right—to aircrafts in order to stay safe and avoid collisions. The system takes the state of the ownship and the relative state of the intruder airplane as inputs, and outputs a score for each of the $5$ possible horizontal maneuvers listed above. Traditionally, this was accomplished using lookup tables but using neural networks has recently become customary.
In their original work, Katz et al. (2017) developed a method called Reluplex for formally verifying that the deep neural networks used for generating the scores of the advisories always satisfy certain “properties” or, as we would call them in our framework, hard constraints. This requires training a model and then verifying that it satisfies certain properties by solving a satisfiability problem. In addition, the original work trains $45$ different models for various values of $\tau$ (the time to collision or loss of vertical separation) and $pra$ (the previously recommended advisory) in order to satisfy strict computational limits imposed by the inference hardware. In our work, we train a single model that generalizes across various values of $\tau$ and $pra$.
In our implementation, we demonstrate a method for learning neural networks that output constrained airplane advisories by construction, rather than engaging in an iterative cycle of training a neural network and separately verifying it until convergence. While some of the properties in the original problem (Katz et al., 2017) are non-convex in nature, we pick all five of the properties that can be encoded in a convex form. The results for this example are presented in Table 5. Additional details for the experiment can be found in Appendix A.6.
Table 5: Results for the HCAS experiments with 17 inequality constraints. Constraint violations are highlighted in red. Standard deviations over 3 runs are shown in parentheses. HardNet-Cvx was warm-started with 250 epochs of soft-penalty training. HardNet-Cvx-NP (no projection) is the same model as HardNet-Cvx but disables the projection at test time. HardNet-Cvx attains feasible solutions with the smallest suboptimality gap among the feasible methods.
@@ 5.4 Airborne Collision Avoidance System (ACAS)
NN 7.98 (1.33) 0.67 (0.26) 0.03 (0.00) 0.11 (0.00) 1.31 (0.03)
NN+Proj 8.15 (1.32) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 6442.21 (250.87)
Soft 7.98 (1.33) 0.67 (0.26) 0.03 (0.00) 0.11 (0.00) 1.33 (0.06)
Soft+Proj 8.15 (1.32) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 6216.16 (332.36)
HardNet-Cvx-NP 7.58 (0.74) 0.38 (0.24) 0.02 (0.00) 0.11 (0.00) 1.66 (0.05)
HardNet-Cvx 7.69 (0.74) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 9367.21 (1383.97)
@@ 6 Conclusion
In this paper, we presented HardNet, a practical framework for constructing neural networks that inherently satisfy input-dependent affine/convex constraints. We proved that imposing these hard constraints does not limit the expressive power of these neural networks by providing universal approximation guarantees. We also demonstrated the utility and versatility of our method across several application scenarios, such as learning solvers for optimization problems, control policies for safety-critical systems, and advisories for aircraft navigation systems. Using HardNet in other application domains that benefit from incorporating domain-specific knowledge is a promising direction for future work. Additionally, we aim to explore developing methods for performing fast projections for problems with more general constraints. Lastly, extending our approach to support other forms of inductive biases, such as equivariances and invariances, would potentially be of great interest.
@@ Acknowledgments
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper. This work was supported in part by MathWorks, the MIT-IBM Watson AI Lab, the MIT-Amazon Science Hub, and the MIT-Google Program for Computing Innovation.
@@ Appendix A Appendix
A.4 Details for the Function Fitting Experiment
@@
A.5 Details for the Safe Control Experiment
@@
where $l$ is is the distance of the body center from the differential drive axis of the unicycle system. However, it is not a valid CBF since the safety condition (12) does not depend on the control input (i.e., $\nabla h_{\text{ellipse}}(x)^{\top}g(x)=0\;\;\forall x$). Instead, we can exploit an higher-order CBF (HOCBF) given by
@@
for some $\kappa>0$. Then, ensuring $h\geq 0$ implies $h\geq 0$ given $h(x(0))\geq 0$, and the safety condition (12) for this $h$ depends on both control inputs $a_{\text{lin}},a_{\text{ang}}$. Refer to Tayal et al. (2024) for detailed explanation.
@@
A.6 Details for the HCAS experiment
A.6.1 Problem Details
In this section, we provide additional details for the HCAS experiment.
@@ A.6.1 Problem Details
      • Constraints on output : $y[4]$ should be the minimal score which translates to
          $y[4]<y[0]$
          $y[4]<y[1]$
          $y[4]<y[2]$
          $y[4]<y[3]$
@@
      • Constraints on output : $y[0]$ should be the minimal score which translates to
          $y[0]<y[1]$
          $y[0]<y[2]$
          $y[0]<y[3]$
          $y[0]<y[4]$
@@
      • Constraints on output : $y[3]$ should be the minimal score which translates to
          $y[3]<y[0]$
          $y[3]<y[1]$
          $y[3]<y[2]$
          $y[3]<y[4]$
@@
      • Constraints on output : $y[0]$ should be the minimal score which translates to
          $y[0]<y[1]$
          $y[0]<y[2]$
          $y[0]<y[3]$
          $y[0]<y[4]$
A.6.2 Training Details
Learning with Warm Start
In addition to the HardNet architecture that consists of a neural network $f_{\theta}$ and a differentiable projection layer $\mathcal{P}$ appended at the end, we propose a training scheme that can potentially result in better-optimized models. For the first $k$ epochs of training, we disable the projection layer and train the plain neural network $f_{\theta}$. Then, from the $(k\!+\!1)$-th epoch, we train on the projected model $\mathcal{P}(f_{\theta})$. During the $k$ epochs of warm start, the neural network $f_{\theta}$ can be trained in a soft-constrained manner by regularizing the violations of constraints. In this paper, we train the HardNet-Aff models without such regularization for simplicity. Incase of HardNet-Cvx we find that performing a warm start is necessary for cvxpylayers since it makes it easier for the convex program solver (SCS) to perform the projection.
Neural network Training Hyperparameters
  • Learning rate (lr) : $1\times 10^{-3}$
  • Number of epochs : $260$

### v2

@@ 5.3 Optimizing Control Policy in Safety-Critical System
NN
421.92 (0.15)
157.62/118.92 (0.90/0.55)
0.49 (0.11) 180.22 (1.25)
Soft
480.10 (0.54)
6.92/3.95 (0.05/0.10)
0.89 (0.76) 180.10 (0.58)
DC3
480.12 (0.59)
6.90/3.93 (0.08/0.13)
24.17 (3.18)
758.50 (61.96)
HardNet-Aff
518.85 (8.71)
0/0 (0/0) 3.16 (0.32) 239.69 (1.46)
In this experiment, we apply HardNet-Aff to enforce safety constraints in control systems. Consider a control-affine system with its known dynamics $f$ and $g$: $\dot{x}(t)=f(x(t))+g(x(t))u(t)$,
Figure 4: Simulated trajectories from a random initial state, with costs shown in parentheses. HardNet-Aff avoids the obstacles while obtaining a low cost value. Even though the soft-constrained method and DC3 appear to avoid obstacles and achieve smaller costs than the other collision-free trajectories, they violate the safety constraints (which are more conservative than hitting the obstacles).
where $x(t)\in\mathbb{R}^{n_{\text{in}}}$ and $u(t)\in\mathbb{R}^{n_{\text{out}}}$ are the system state and the control input at time $t$, respectively. For safety reasons (e.g., avoiding obstacles), the system requires $x(t)\in\mathcal{X}_{\text{safe}}\subset\mathbb{R}^{n_{\text{in}}}$ for all $t$. We translate this safety condition into a state-dependent affine constraint on the control input using a control barrier function (CBF) $h:\mathbb{R}^{n_{\text{in}}}\rightarrow\mathbb{R}$ [4]. Suppose its super-level set $\{x\in\mathbb{R}^{n_{\text{in}}}|h(x)\geq 0\}\subset\mathcal{X}_{\text{safe}}$ and $h(x(0))\geq 0$. Then, we can ensure $h(x(t))\geq 0\;\forall t\geq 0$ by guaranteeing
@@
We consider controlling a unicycle system to minimize the cost over trajectories while avoiding collisions with two elliptical obstacles, each presented with a CBF (see Appendix A.7 for details). Then, we can formulate the problem as an optimization problem with its objective function being the expected cost over the trajectory $x(t)$ generated by a parameterized state feedback policy $\pi_{\theta}:\mathbb{R}^{n_{\text{in}}}\rightarrow\mathbb{R}^{n_{\text{out}}}$ from random initial point $x(0)\sim\mathcal{D}$:
@@
where $Q$ is the state cost matrix and $R$ is the control cost matrix.
Given a nominal controller $\pi_{\text{nom}}:\mathbb{R}^{n_{\text{in}}}\rightarrow\mathbb{R}^{n_{\text{out}}}$ designed without obstacle considerations, a conventional approach to find a safe controller is to solve the following quadratic program at each $x(t)$:
@@
The downside of this method is that the controller cannot optimize a cost/reward over trajectories, as it only stays close to the nominal controller. Instead, we can do so by training neural network policies $\pi_{\theta}(x):=\pi_{\text{nom}}(x)+f_{\theta}(x)$ with neural networks $f_{\theta}$. For computation, we approximate (11) by minimizing the costs of rolled-out trajectories from randomly sampled initial states. As shown in Fig. 4 and Table 4, HardNet-Aff consistently guarantees safe trajectories with low costs.
@@ 6 Conclusion
In this paper, we presented HardNet, a practical framework for constructing neural networks that inherently satisfy input-dependent affine/convex constraints. We proved that imposing these hard constraints does not limit the expressive power of these neural networks by providing universal approximation guarantees. We demonstrated the utility and versatility of our method across several applications, such as learning with piecewise constraints, learning optimization solvers, and optimizing control policies in safety-critical systems. Using HardNet in other application domains that benefit from incorporating domain-specific knowledge is a promising direction for future work. One such example is in [59]. Also, we aim to explore developing methods for performing fast projections for problems with more general constraints. Lastly, extending our approach to support other forms of inductive biases, such as equivariances and invariances, would potentially be of great interest.
@@ Acknowledgments
The authors thank Anoopkumar Sonar for discussions and helpful input on an early version of this manuscript. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper. This work was supported in part by MathWorks, the MIT-IBM Watson AI Lab, the MIT-Amazon Science Hub, and the MIT-Google Program for Computing Innovation.
@@ Appendix A Appendix
Figure 5: Visualization of 100 gradient descent steps for training a HardNet-Aff model on a single datapoint (first row) and two datapoints (second row) from the same initialization, using two different learning rates (0.01 and 0.1). With the smaller learning rate, training on a single datapoint results in a zero gradient due to the enforcement layer (top left). However, when training on both datapoints, the vanishing gradient for the first datapoint is mitigated by the nonzero gradient from the second datapoint (bottom left). Also, using the larger learning rate enables the model to avoid the vanishing gradient issue, even when trained on the single datapoint (top right).
A.4 Gradient Properties of HardNet-Aff
This section investigates how the enforcement layer in HardNet-Aff affects gradient computation. For simplicity, we focus on the case where $n_{\text{eq}}=0$ (i.e., no equality constraints). For a datapoint $(x,y)$, consider the loss function $\ell\big{(}\mathcal{P}(f_{\theta}(x)),y\big{)}$. Using the chain rule, the gradient is given by:
$\nabla_{\theta}\ell\big{(}\mathcal{P}(f_{\theta}(x)),y\big{)}^{\top}=\frac{\partial\ell\big{(}\mathcal{P}(f_{\theta}(x)),y\big{)}}{\partial\mathcal{P}(f_{\theta}(x))}\frac{\partial\mathcal{P}(f_{\theta}(x))}{\partial f_{\theta}(x)}\frac{\partial f_{\theta}(x)}{\partial\theta}.$ (37)
Here, the Jacobian of the enforcement layer $\frac{\partial\mathcal{P}(f_{\theta}(x))}{\partial f_{\theta}(x)}\in\mathbb{R}^{m\times m}$ plays a key role. Let $v_{i}:=\mathds{1}\{a_{i}(x)^{\top}f_{\theta}(x)>b_{(i)}(x)\}$ indicate whether the $i$-th constraint is violated by $f_{\theta}(x)$. Then,
$\frac{\partial\mathcal{P}(f_{\theta}(x))}{\partial f_{\theta}(x)}=I-A^{+}\begin{bmatrix}v_{1}a_{1}^{\top}\\ \vdots\\ v_{n_{\text{ineq}}}a_{n_{\text{ineq}}}^{\top}\end{bmatrix}.$ (38)
Two critical properties of this Jacobian can lead to zero gradient in (37). First, if the number of constraints equals the output dimension ($n_{\text{ineq}}=m$) and all constraints are violated ($v_{i}=1\;\forall i$), then the Jacobian becomes zero, causing the gradient (37) to vanish. Note that this issue never happens when $n_{\text{ineq}}<m$. Second, for each $i\in\{1,2,\dots,n_{\text{ineq}}\}$, the following holds:
$a_{i}^{\top}\frac{\partial\mathcal{P}(f_{\theta}(x))}{\partial f_{\theta}(x)}=a_{i}^{\top}-a_{i}^{\top}A^{+}\begin{bmatrix}v_{1}a_{1}^{\top}\\ \vdots\\ v_{n_{\text{ineq}}}a_{n_{\text{ineq}}}^{\top}\end{bmatrix}=a_{i}^{\top}-e_{i}^{\top}\begin{bmatrix}v_{1}a_{1}^{\top}\\ \vdots\\ v_{n_{\text{ineq}}}a_{n_{\text{ineq}}}^{\top}\end{bmatrix}=a_{i}^{\top}-v_{i}a_{i}^{\top}.$ (39)
This implies that if the gradient of the loss with respect to the projected output, $\big{(}\frac{\partial\ell\big{(}\mathcal{P}(f_{\theta}(x)),y\big{)}}{\partial\mathcal{P}(f_{\theta}(x))}\big{)}^{\top}\in\mathbb{R}^{m}$, lies in the span of $\{a_{i}|i\in\{1,\dots,n_{\text{ineq}}\},v_{i}=1\}$, then the overall gradient (37) becomes zero. This case in fact subsumes the first case, as when $n_{\text{ineq}}=m$ and $v_{i}=1\;\forall i$, the constraint vectors set spans the entire output space $\mathbb{R}^{m}$. However, such cases are infrequent in practical settings, particularly when the model is trained on batched data. Even when zero gradients occur for certain datapoints, they can be offset by nonzero gradients from the other datapoints within the batch. This averaging effect allows the model to update in a direction that decreases the overall loss function, mitigating the issue effectively.
We demonstrate the gradient behaviors discussed earlier using simple simulations of training HardNet-Aff with the conventional gradient-descent algorithm. Consider two datapoints: $d_{1}=(-1,[-0.5,0.5]^{\top})$ and $d_{2}=(1,[0.5,0]^{\top})$, and a neural network $f_{\theta}:\mathbb{R}\rightarrow\mathbb{R}^{2}$ with two hidden layers, each containing 10 neurons with $\mathsf{ReLU}{}$ activations. The model enforces the input-independent constraint $[0,1]\mathcal{P}(f_{\theta})(x)\leq 0.9$ using HardNet-Aff. Starting from the same initialization, the model is trained to minimize the squared error loss, first on $d_{1}$ alone and then on both datapoints, using two different learning rates (0.01 and 0.1), as shown in Fig. 5.
Initially, $f_{\theta}$ violates the constraint on both datapoints. When trained on $d_{1}$ alone with a learning rate of 0.01, the optimization path converges to a point where the gradient of the loss w.r.t. the projected output is orthogonal to the gradient boundary, causing the overall gradient (37) to vanish. However, when the model is trained on both datapoints, the vanishing gradient for $d_{1}$ is mitigated by the nonzero gradient for $d_{2}$, enabling the model to achieve target values strictly within the feasible set. Furthermore, using a larger learning rate (0.1) allows the model to avoid the vanishing gradient issue and reach the target value even when trained solely on $d_{1}$.
A.5 Learning with Warm Start
In addition to the HardNet architecture shown in Figure 1 that consists of a neural network $f_{\theta}$ and a differentiable enforcement layer with a projection $\mathcal{P}$ appended at the end, we propose a training scheme that can potentially result in better-optimized models. For the first $k$ epochs of training, we disable the enforcement layer and train the plain neural network $f_{\theta}$. Then, from the $(k\!+\!1)$-th epoch, we train on the projected model $\mathcal{P}(f_{\theta})$. During the $k$ epochs of warm start, the neural network $f_{\theta}$ can be trained in a soft-constrained manner by regularizing the violations of constraints. In this paper, we train the HardNet-Aff models without the warm-start scheme for simplicity.
A.6 Details for the Learning with Piecewise Constraints Experiment
@@
The results in Section 5.1 show HardNet-Aff can help generalization on unseen regimes by enforcing constraints. In this section, we provide additional results that train the models on data spanning the entire domain of interest $[-2,2]$. As shown in Figure 6 and Table 5, the models exhibit similar generalization performances while HardNet-Aff satisfy the constraints throughout the training.
Figure 6: Learned functions at the initial (left) and final (right) epochs for the function fitting experiment. The models are trained on the samples indicated with circles, with their MSE distances from the true function shown in parentheses. HardNet-Aff adheres to the constraints from the start of the training. On the other hand, the baselines violate the constraints throughout the training.
Table 5: Results for the function fitting experiment. HardNet-Aff attains a comparable MSE distance from the true function as other methods without any constraint violation. The max, mean, and the number of constraint violations are computed out of 401 test samples. Violations are highlighted in red. Standard deviations over 5 runs are shown in parentheses.
[…]
NN 0.04 (0.01) 0.73 (0.07) 0.02 (0.01) 42.40 (5.31) 1.54 (0.02) 2.45 (0.80)
Soft 0.05 (0.01) 0.69 (0.04) 0.02 (0.00) 28.60 (2.42) 1.82 (0.37) 2.54 (0.87)
DC3 0.04 (0.01) 0.48 (0.03) 0.01 (0.00) 30.60 (1.85) 50.99 (0.11) 14.43 (0.99)
HardNet-Aff 0.06 (0.01) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 7.04 (0.23) 4.85 (0.78)
A.7 Details for the Safe Control Experiment
@@
where $l$ is the distance of the body center from the differential drive axis of the unicycle system. However, it is not a valid CBF since the safety condition (10) does not depend on the control input (i.e., $\nabla h_{\text{ellipse}}(x)^{\top}g(x)=0\;\;\forall x$). Instead, we can exploit a higher-order CBF (HOCBF) given by
@@
for some $\kappa>0$. Then, ensuring $h\geq 0$ implies $h\geq 0$ given $h(x(0))\geq 0$, and the safety condition (10) for this $h$ depends on both control inputs $a_{\text{lin}},a_{\text{ang}}$. Refer to Tayal et al. [61] for a detailed explanation.
@@
A.8 Additional Scenario: Airborne Collision Avoidance System (ACAS)
In this scenario, we consider learning an airborne collision avoidance system for unmanned aircraft (ACAS Xu) used by Katz et al. [32]. The goal of the system is to provide horizontal maneuver advisories—clear of conflict, weak left, strong left, weak right, strong right—to aircrafts in order to stay safe and avoid collisions. The system takes the state of the ownship and the relative state of the intruder airplane as inputs, and outputs a score for each of the $5$ possible advisories listed above. Traditionally, this was accomplished using lookup tables but using neural networks has recently become customary.
In their original work, Katz et al. [32] developed a method called Reluplex for formally verifying that the deep neural networks used for generating the scores of the advisories always satisfy certain hard constraints. This requires training a model and then verifying that it satisfies the constraints by solving a satisfiability problem. Instead, we can learn neural networks that output constrained airplane advisories by construction, rather than engaging in an iterative cycle of training a neural network and separately verifying it until convergence. While some of the properties in the original problem [32] are non-convex in nature, we introduce five of the properties that can be encoded in a convex form below.
A.8.1 Problem Details
@@
      • Constraints on output : $y[4]$ should be the minimal score which translates to $y[4]\leq\min\{y[0],y[1],y[2],y[3]\}$
@@
      • Constraints on output : $y[0]$ should be the minimal score which translates to $y[0]\leq\min\{y[1],y[2],y[3],y[4]\}$
@@
      • Constraints on output : $y[3]$ should be the minimal score which translates to $y[3]\leq\min\{y[0],y[1],y[2],y[4]\}$
@@
      • Constraints on output : $y[0]$ should be the minimal score which translates to $y[0]\leq\min\{y[1],y[2],y[3],y[4]\}$
A.9 Related Work in Neuro-Symbolic AI
HardNet also aligns with the objectives of Neuro-symbolic AI, a field that has gained significant attention in recent years for its ability to integrate complex background knowledge into deep learning models. Unlike HardNet, which focuses on algebraic constraints, the neuro-symbolic AI literature primarily addresses logical constraints. A common approach in this field is to softly impose constraints during training by introducing penalty terms into the loss function to discourage constraint violations [68, 24, 7, 57]. While these methods are straightforward to implement, they do not guarantee constraint satisfaction. In contrast, works such as Giunchiglia & Lukasiewicz [26], Ahmed et al. [2], Giunchiglia et al. [27] ensure constraints are satisfied by embedding them into the predictive layer, thus guaranteeing compliance by construction. Another line of research maps neural network outputs into logical predicates, ensuring constraint satisfaction through reasoning on these predicates [42, 50, 64].